OpenStudy (anonymous):

If a, b and c are odd integers, which of the following expressions must be an even integer? a + b + c a(b + c) ab + bc + ca a(b + c - 1)

OpenStudy (mepitt):

i think c

OpenStudy (whovianchick):

No, it's B. plug in odd numbers for all of the variables, and you get B every time

OpenStudy (anonymous):

First fact needed: ODD + ODD = EVEN This is always true. And, ODD + EVEN = ODD. So, a + b + c is (ODD + ODD + ODD) = (ODD + ODD) + ODD = EVEN + ODD = ODD.

OpenStudy (anonymous):

@KaylaRdz0405

OpenStudy (anonymous):

Now, we need more facts. (EVEN)(EVEN) = EVEN and (EVEN)(ODD) = EVEN; that is, multiplying by an even number always given you an even number. Do you see how that can help you with a(b + c)?

OpenStudy (anonymous):

@TINY10207

OpenStudy (anonymous):

I'll get you started: ODD(ODD + ODD) = ?

OpenStudy (anonymous):

EVEN

OpenStudy (anonymous):

RIght... ODD(ODD + ODD) = ODD(EVEN) = EVEN. Yep!

OpenStudy (whovianchick):

It's B, I've already answered. Plug in numbers, and you'll see I'm right

OpenStudy (anonymous):

Now, how about the others? (Perhaps there is more than one that must always be even...)

OpenStudy (kaylardz0405):

i agree with whovianchick

OpenStudy (anonymous):

Plugging in #s is a good approach. Sometimes it's useful to use some other ideas, too.

OpenStudy (whovianchick):

@OneMathCat The problem is already solved. Stop rambling.

OpenStudy (anonymous):

Yep!! I do have a tendency to ramble ... sorry! Have a good day!

OpenStudy (anonymous):

THANK U @whovianchick

OpenStudy (whovianchick):

no prob :)